OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 slope = -5; y-intercept = (0, -6) slope = five thirds; y-intercept = (0, 2) slope = -three fifths; y-intercept = (0, 2) slope = five thirds; y-intercept = (0, -6)

OpenStudy (anonymous):

You need to change the function into the form y=mx+b where m is the slope. Set the 0 to x and you will have the y intercept.

OpenStudy (anonymous):

Well you will want to make the formula into y-intercept form which is y=mx+b. m is the slope and b is the y-intercept

OpenStudy (anonymous):

To change into y-intercewpt form you will first subtract 5x from both sides to get -3y = -6 - 5x

OpenStudy (anonymous):

so it would be c?

OpenStudy (anonymous):

Next you will divide -3 from both sides to get y=2-5/3x

OpenStudy (anonymous):

Finally you will fix into y-intercept form which will be y=5/3x+2

OpenStudy (anonymous):

It would be B

OpenStudy (anonymous):

oh okay i c it now

OpenStudy (anonymous):

another?

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (2, -3) and is perpendicular to y = two thirdsx + 3. y = -three halvesx y = three halvesx - 6 y = -two thirdsx - 5 thirds y = two thirdsx - 13 thirds

OpenStudy (anonymous):

If you need another question answer post it as a new question @JOshua_Lewis